OpenStudy (pottersheep):

Grade 12 math help please!! For what value of b will the polynomial p(x) = -2x^3 + bx^2 - 5x +2 have the same remainder when it is divided by x - 2 and x+1?

OpenStudy (anonymous):

okay I will show you:

OpenStudy (pottersheep):

thank you so much!

OpenStudy (pottersheep):

:(

OpenStudy (anonymous):

thinking

OpenStudy (pottersheep):

oh ok sorry

OpenStudy (anonymous):

sorry, it too late for me I will hand this over to lonesome else?

OpenStudy (pottersheep):

ok

OpenStudy (pottersheep):

thanks anyways

OpenStudy (callisto):

Here it goes~ Have you learnt remainder theorem yet?

OpenStudy (callisto):

@pottersheep Are you there??

OpenStudy (pottersheep):

Yes sorry, we have!

OpenStudy (callisto):

So, let R be the remainder when p(x) is divided by x-2. By remainder theorem, what would you get? I mean p(_) = ...?

OpenStudy (pottersheep):

ok one sec

OpenStudy (pottersheep):

So i'd sub in 2 for x

OpenStudy (callisto):

Yup!

OpenStudy (pottersheep):

3?

OpenStudy (pottersheep):

sorryryyryryryr

OpenStudy (pottersheep):

Sorry what do I do with the b though?

OpenStudy (pottersheep):

4b -24

OpenStudy (callisto):

Right, so for p(2), you get p(2) = 4b - 24 = R, agree? (By remainder theorem, p(a) = R , when p(x) is divided by x-a)

OpenStudy (pottersheep):

Right, okay

OpenStudy (callisto):

And what is the remainder when p(x) is divided by x+1 ?

OpenStudy (pottersheep):

b - 5

OpenStudy (callisto):

Not really.. Remainder = p(-1) = -2(-1)^3 + b(-1)^2 - 5(-1) +2 = ...?

OpenStudy (pottersheep):

Opps, sorry! I subbed in 1!

OpenStudy (pottersheep):

9??

OpenStudy (pottersheep):

b + 9!

OpenStudy (callisto):

Yes. Note that the remainder is p(x) = R when p(x) is divided by x-a. Now, p(x) is divided by x+1, so a = -1 in the case. [ x-(-1) = x+1]

OpenStudy (pottersheep):

True, sorry!

OpenStudy (callisto):

From the question, the polynomial p(x) = -2x^3 + bx^2 - 5x +2 have the same remainder when it is divided by x - 2 and x+1. So, equate the two remainder you got and solve b. Got it?

OpenStudy (callisto):

Oh.. It's okay!

OpenStudy (pottersheep):

Ohhhhhhhhh, so make them equal to each other?? Yes I get it, thank you!!!!!!

OpenStudy (callisto):

Yes. Since they have the SAME remainder! You're welcome :)